# X-GAN: A Generative AI-Powered Unsupervised Model for High-Precision Segmentation of Retinal Main Vessels toward Early Detection of Glaucoma

### v1

@@ 4.1.1 GSS-RetVein
GSS-RetVein OCTA-500 ROSE
@@
X-GAN (MAX) 99.89 99.94 99.66 99.83 99.73 99.86 99.80 99.89 99.57 99.78
Table 1: Comparison X-GAN With other models on GSS-RetVein, OCTA-500 and ROSE. ($\times$100%)
Data Collection and Quality Control: our partner medical institutions’s institutional review board (IRB) approved this study, which followed the principles of the Declaration of Helsinki. Since the study was retrospective, the IRB waived the requirement for informed consent from patients.
4.1.2 OCTA-500
4.1.3 ROSE
4.2 Implement
4.2.1 Hyperparameter
For X-GAN, its generator adopts a ResNet 9-block architecture, while the discriminator utilizes PatchGAN (70$\times$70). The optimizer is Adam with a learning rate of $2\times 10^{-4}$ and momentum parameters $\beta_{1}=0.5$, $\beta_{2}=0.999$. X-GAN is trained for 50 epochs, with a linear learning rate decay to 0 after 25 epochs, with evaluation performed using 10-fold cross-validation. Data augmentation includes: random rotations ($k\times 90^{\circ}\pm 10^{\circ}$), flipping and contrast adjustments.
The hardware specifications for training and testing include 2 Tesla V100 GPUs (2 $\times$ 32GB), 64GB of RAM, 8 CPU cores per node, and a total of 6 nodes.
4.2.2 Evaluation Metric
We use Intersection over Union (IoU), also known as the Jaccard Index, to measure similarity between the predicted segmentation and the ground truth: $\text{IoU}=|P\cap G|/|P\cup G|$. This metric ranges from 0 to 1, with 1 indicating perfect segmentation. Additionally, we use the Dice Coefficient, another metric for segmentation accuracy that calculates overlap between the prediction and ground truth: $\text{Dice}=2\times|P\cap G|/(|P|+|G|)$. In both formulas, $P$ represents predicted vessel pixels, $G$ the ground truth vessel pixels, $|P\cap G|$ the count of correctly predicted vessel pixels, and $|P\cup G|$ the combined set of ground truth and predicted vessel pixels. All in all, IoU is used to evaluate segmentation accuracy of the main vessels, while the Dice Coefficient assesses accuracy at finer vessel ends relative to the main vessels. For vascular structural enhancement, we use SSIM (Structural Similarity Index) and MSE (Mean Squared Error). SSIM (range: -1 to 1) assesses similarity based on luminance, contrast, and structure, while MSE quantifies pixel-wise differences, with lower values indicating higher similarity. While SSIM aligns with human perception, PSNR and MSE focus on absolute pixel differences.
(a) Comparison of models, trained on GSS-RetVein and OCTA and tested on ROSE.
(b) Comparison of models, trained on ROSE and tested on GSS-RetVein and OCTA.
(c) Comparison of models, trained on GSS-RetVein or OCTA and tested on each other.
Figure 6: Conducting mixed testing on different models using mixed datasets. ($\times$100%)
5 Experimental Result
@@
GSS-RetVein OCTA-500 ROSE
@@
Table 2: Comparison vessel structure refinement images with real OCTA images. ($\times$100%)
5.1 Comparative Experiment
As presented in Tab. 1, experimental findings demonstrate that while models such as MedSAM, U-Net++, and YOLOv11-x exhibit competitive performance, X-GAN consistently outperforms all competing models across varying dataset resolutions, achieving the highest scores and excelling in retinal vessel segmentation, where both of the evaluation matrixs are even close to 100%, highlighting its superior segmentation accuracy and underscoring its robustness in fine-grained vascular structures segmentation.
As shown in Fig. 6(a), given the comparable sizes of GSS-RetVein and OCTA-500 (retinal vessel subset) and the smaller size of ROSE, we trained models on the first two and tested on the latter. X-GAN consistently achieved the highest IoU and Dice scores on ROSE, outperforming S2VNet, CauSSL, and MedSAM. To further assess generalization, we trained models on ROSE and tested them on GSS-RetVein and OCTA-500 (Fig. 6(b)), where X-GAN again achieved the highest scores, followed by MedSAM and CauSSL. Finally, Fig. 6(c) evaluates models across different scanning resolutions, with X-GAN consistently leading in segmentation accuracy. While MedSAM and UniverSeg performed well, U-Net and its variants showed strong but relatively lower performance. These results highlight GSS-RetVein’s high quality and X-GAN’s superior segmentation capability.
Note: the highest indicator is shown in bold, while the second highest is marked with an underline. We do not consider X-GAN (MAX) in all Tabs. 1, 6 and 3.
5.2 Ablation Experiment
For baseline GAN model, GAN [13] and CycleGAN [66] are selected. For baseline segmentor, U-Net [49], U-Net++ [64], Attention U-Net [44] and MedSAM [39] are selected. For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Tab. 3, since main blood vessels make up a small proportion, the value calculated is also relatively low. When the minimum radius ratio is found, like X-GAN (MAx), aligning with biostatistical properties [22, 12, 5, 43, 7]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
@@ 5.2 Ablation Experiment
Figure 7: Ablation experiments of key parameters. ($\times$100%)
Also, as shown in Eq. 1 and Fig. 7(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of X-GAN and X-GAN (MAX) are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar. Consequently, their curves largely overlap, with X-GAN (MAX) exhibiting a slight increase.
5.3 Evaluation of Vessel Structure Refinement
The comparison between CycleGAN [66] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%. CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets. The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
Indicator SLSMs DFS (ours)
GPU $\checkmark$ $\times$
Annotation $\checkmark$ $\times$
Training $\checkmark$ $\times$
Explainability $\times$ $\checkmark$
2D and 3D $\times$ $\checkmark$
Generalization $\times$ -
Table 4: Comparison of performance indicators between DFS (our) and SLSMs. ($\times$100%) (- indicates strong performance for a specific class.)
(a) Segmentation of Main Retinal Vessels (2D)
(b) Original Image (2D)
(c) Original Image (3D)
(d) Segmentation (2D)
(e) Segmentation (3D)
Figure 8: Segmentation results of X-GAN on GSS-RetVein (2D and 3D).
@@ 5.4 Efficiency Evaluation
All in all, as shown in Tab. 4, compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
What is more, in more detail, the advantages of our approach are as follows in Tab. 5, our segmentor has demonstrated unsurpassed superiority.
@@
Table 5: Comparison of performance details between DFS (our) and SLSMs.
@@ 5.5 Visualization
As shown in Fig. 8, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 8(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 8(e) preserving vessel integrity and offering a clearer structure than the original Fig. 8(c).
@@ 5.6 Analysis
Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Tabs. 1, 6 and 3). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [22, 12, 5, 43, 7]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results. Studies [34, 56] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models. There are many examples: Faster R-CNN [48] and Mask R-CNN [16], YOLOv3 [47], YOLOv4 [4] and YOLOv5 [17], DeepSeek [14, 33] and ChatGPT [1].
To the best of our knowledge, X-GAN is the best model currently, tailored to segment OCTA retinal main vessel.
@@ 6 Conclusion
In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. We also introduce GSS-RetVein, a high-resolution dataset featuring both 2D and 3D retinal vascular structures, offering sharper vessel boundaries and detailed capillary networks compared to existing datasets, establishing it as a superior benchmark for vessel segmentation. Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
Acknowledgements
This research is sponsored by NASA 80NSSC22K0144, UTSW GMO No. 241213, and NIH 1R01AG083179-01.

### v3

@@ 5.1 Comparative Experiment
UTSW OCTA-500 ROSE
@@
As shown in Table 1, experimental results indicate that while MedSAM, U-Net++, and YOLOv11-x perform competitively, X-GAN consistently outperforms them across resolutions, achieving near-perfect evaluation metrics and excelling in fine-grained retinal vessel segmentation.
As shown in Table 2, comparison between CycleGAN [35] and CycleGAN (ours revised version) shows a high structural similarity, with SSIM values ranging from 91.01% to 96.04%. CycleGAN (ours) consistently outperforms CycleGAN, achieving higher SSIM and lower MSE across all datasets. The small SSIM difference (less than 4.31%) indicates that both models generate highly similar images, but CycleGAN (ours) preserves image structure better and reduces pixel-level errors, making it the superior model.
Table 2: Comparison Vessel Structure Refinement with Real OCTA ($\times$100%)
@@
UTSW OCTA-500 ROSE
@@
5.2 Ablation Eperiment
For baseline GAN model, GAN [34] and CycleGAN [35] are selected. For baseline segmentor, U-Net [10], U-Net++ [7], Attention U-Net [12] and MedSAM [8] are selected. For X-GAN (CycleGAN (ours) + DFS), the dirrerent values of $R_{min}$ are tested to evaluate the model segmentation accuracy. As shown in Table 3, since main blood vessels make up a small proportion, the value calculated is also relatively low, aligning with biostatistical properties [30, 31, 33, 32]—segmentation accuracy peaks, reinforcing the model’s strong clinical interpretability.
Table 3: Ablation Experiment of X-GAN ($\times$100%)
@@ 5.2 Ablation Eperiment
Figure 3: Ablation Experiments of Key Parameters ($\times$100%)
Also, as shown in Fig. 3(b), a greater number of generated nodes $N$ leads to smoother vessel edges, a more complete vascular structure, and improved segmentation accuracy. Given that the R values of X-GAN are relatively low, their practical significance indicates minimal differences in the segmented vessels, as the number of nodes remains similar.
@@ 5.3 Efficiency Evaluation
Compared to SLSMs, our segmentor offers these advantages: faster inference (DFS does not need GPU and training, but SLSMs need GPU for training and testing), eliminates the need for large-scale annotations, and ensures full preservation of primary vessel integrity. It features an adjustable capillary filtering parameter $R_{min}$, maintains strong generalization across all vessel OCTA datasets without retraining, and is directly applicable to 3D extensions without additional design.
Table 4: Comparison of Performance Details between DFS (our) and SLSMs
@@
Compared to the segmentor, our adjusted DFS offers unmatched advantages, as shown in Table 4. It requires no separate training or labeling, as DFS is a parameter-efficient algorithm. By eliminating manual annotation, it removes subjective bias from doctors. Additionally, its radius-based segmentation avoids pixel mapping errors in SLSMs, such as interference at vessel edges, inaccuracies at junctions, and mask coverage errors. What is more, the key issue is that both SLSMs and GAN+SLSMs frameworks are susceptible to fitting problems in segmentation tasks. This arises from the training limitations of the SLSMs model and the nature of the data. In contrast, DFS, as an algorithm rather than a model, is unaffected by this issue.
@@ 5.4 Visualization
As shown in Fig. 4, our 2D and 3D segmentation results exhibit smooth, well-defined vessel curves, confirming our prior analysis. In Fig. 4(a), the segmented vessels are exceptionally smooth, free of pixel artifacts or overflow. DFS effectively filters out vessels below the threshold, eliminating noise and misclassified capillaries. Extending this to 3D, we achieve precise main vessel extraction, with Fig. 4(c) preserving vessel integrity and offering a clearer structure than the original image (Fig. 4(b)).
(a) Segmentation of Main Retinal Vessels (2D)
(b) Original Image (3D)
(c) Segmentation (3D)
Figure 4: Segmentation of X-GAN (2D and 3D)
@@ 5.5 Analysis
Optimal selection of $R_{min}$ values allows X-GAN to achieve exceptionally high segmentation accuracy, outperforming the second-best model by nearly 3 percentage points, nearly 100% (Table 1 and Table 3). Unlike many SLSMs that have been magically tuned to have very poor interpretability, $R_{min}$ is based on biostatistics of glacucoma [30, 31, 33, 32]. We only knew the biological information of the main blood vessel radius and the form of its data expression, and just adjusted X-GAN to achieve amazing segmentation results. Unlike GAN+SLSMs or SLSMs, X-GAN does not require additional data to avoid fitting issues. Its segmentor is efficient, computationally friendly, and free from pixel mapping errors associated with SLSMs segmentation. Studies [28, 29] have also shown the similar idea: with a deep understanding of data characteristics, the right training strategies, and optimized parameters, even CNNs can outperform transformer-based models.
@@ 6 Conclusion
In this paper, we propose X-GAN, an unsupervised model for ultra-high-precision segmentation of OCTA retinal main vessels, integrating biostatistical vessel radius properties with a GAN-enhanced DFS algorithm to achieve near-perfect accuracy without labeled data, training, or high-performance GPUs. Extensive experiments demonstrate that X-GAN consistently surpasses SOTA models, achieving nearly 100% segmentation accuracy, confirming its robustness and generalizability. Our DFS-based segmentation effectively isolates main vessels while filtering out capillary noise, ensuring strong biological interpretability and clinical relevance.
7 Future Work
With access to state-of-the-art glaucoma detection equipment, we have curated a high-quality image dataset featuring precisely annotated retinal vascular images. This dataset is well-suited for glaucoma pathology research and computer vision.